# EquiBot: SIM(3)-Equivariant Diffusion Policy for Generalizable and Data Efficient Learning

### v1

@@ Abstract
Building effective imitation learning methods that enable robots to learn from limited data and still generalize across diverse real-world environments is a long-standing problem in robot learning. We propose EquiBot, a robust, data-efficient, and generalizable approach for robot manipulation task learning. Our approach combines SIM(3)-equivariant neural network architectures with diffusion models. This ensures that our learned policies are invariant to changes in scale, rotation, and translation, enhancing their applicability to unseen environments while retaining the benefits of diffusion-based policy learning, such as multi-modality and robustness. We show on a suite of 6 simulation tasks that our proposed method reduces the data requirements and improves generalization to novel scenarios. In the real world, with 10 variations of 6 mobile manipulation tasks, we show that our method can easily generalize to novel objects and scenes after learning from just 5 minutes of human demonstrations in each task. Website: https://equi-bot.github.io/
@@ 1 Introduction
The quest for fully autonomous robotic agents has often stumbled over the unpredictability and variability of real-world environments. While many existing visuomotor policies trained via imitation learning are effective in controlled settings, they require a significant amount of data and even then struggle to adapt when evaluated with unfamiliar objects in unfamiliar environments. These limitations restrict the practical deployment of autonomous robots into the real world and amplify the need for policies that can learn from limited data, yet perform robustly across diverse scenarios.
In this work, we introduce EquiBot, an equivariant policy learning architecture based on diffusion models. With an equivariant neural network architecture, the model output is guaranteed to scale, translate, and rotate with the inputs, even if the model is not fully trained. This allows a learned policy to generalize to unseen scenarios that depart drastically from the scenarios in which it is trained. The equivariant architecture also brings data efficiency benefits, since an equivariant policy can infer how to react to object placements and poses that are in distribution but not sufficiently demonstrated, thus bringing better performance than a non-equivariant policy. Using a policy architecture based on diffusion models also offers robust policy learning performance, including support for idle actions and multi-modal behaviors. With our policy design, EquiBot can enable a wide range of household manipulation tasks from just a handful of single-view human demonstration videos. At test time, our method takes single-view scene point clouds and robot proprioception information as input, and outputs sequences of robot end-effector velocity actions as outputs.
We evaluate our method in both simulation and real-world experiments. In simulation, we evaluate on a suite of 6 tasks (box closing, cloth folding, object covering, push T, can pick-and-place, and square nut assembly). We show that compared to prior work in imitation learning and equivariant visuomotor policy learning, our method is both more data-efficient and generalizes better in unseen scenarios. In real robot experiments, we quantitatively test our method in 6 real-world mobile manipulation tasks in real bedroom and living room settings. Our experiments cover a wide range of everyday tasks, including pushing a chair towards a desk, closing the door of a laundry machine, folding towels, making the bed, and closing a suitcase. These tasks range from single-robot pick-and-place tasks to multi-robot tasks involving deformable and articulated objects. We show that our method can successfully perform the learned tasks with unseen objects and scenes from just 5 minutes of human demonstrations, outperforming competing baselines that either rely on augmentations to achieve generalization or do not utilize a diffusion process to predict actions.
@@ 2 Related Work
Data-efficient imitation learning. Recent imitation learning approaches assume large quantities of data to be available for learning manipulation policies [1, 2]. To reduce the amount of data that policy training requires per task, some works formulate a one-shot or few-shot imitation learning setup, where a policy is trained on multi-task demonstration data and can then output actions for a novel task after seeing one or more demonstrations [3, 4, 5, 6]. While this approach improves the data efficiency of policy learning per task, it requires the availability of multi-task data. Other prior works achieve imitation learning from small demonstration datasets with active learning methods, such as Bayesian optimization, but these methods are often limited to small action spaces and open-loop settings [7]. In contrast to prior works, we show that embedding equivariance into the policy architecture can effectively improve the data efficiency of the imitation learning algorithm, allowing a robust manipulation policy to be learned with just a handful of demonstrations.
@@
Equivariance in robot manipulation. To help learned robot policies generalize to unseen environments and object placements, prior works [8, 9, 10] use data augmentation to improve cross-domain transfer of the learned policy. However, these methods increase training time significantly and do not guarantee generalization to visual appearances, object scales, and poses that are unseen in the training data distribution even after the augmentation. In contrast, utilizing equivariant representations in policy learning allows the learned policies to generalize to objects and initial conditions not previously seen at training time. Prior works have explored the use of equivariance in robot manipulation in several different setups [11, 12, 13, 14, 15, 16, 17, 18, 19, 20], but most of them either focus on only simple pick-and-place like tasks, do not support closed-loop policies, or do not support scale equivariance. Closely related to our work, [21] developed a SIM(3)-equivariant visuomotor policy learning method that can handle non-pick-and-place tasks with deformable and articulated objects. However, this method demonstrated support only for simple behavior cloning. Our method combines SIM(3)-equivariant neural network architectures with diffusion policy [1], thus can be robustly trained and generalizes to unseen object appearance, initial states, scales, and poses.
@@
Equivariant diffusion architectures. Prior works have integrated equivariance in diffusion models in various non-robotics domains, including molecule generation [22, 23, 24] and drug design [25]. Some works have attempted to integrate equivariant architectures in diffusion models for robotics [26, 27, 28]. EquiDiff [26] combines equivariance with diffusion policy, but only handles SO(2)-equivariance with simple 2D trajectories, which is insufficient for most 3D manipulation tasks. Diffusion-EDFs [27] take point cloud observations as input and output a single target end-effector pose. This formulation makes the work only applicable to pick-and-place tasks. EDGI [28] proposed an open-loop policy and showed it in simple 2D tasks only. Compared to prior works, our work proposes an equivariant diffusion policy that supports closed-loop 3D manipulation tasks.
Please see supplementary materials (Section A) for further discussion of related work, including the relationship between our work and prior work using diffusion policy with point cloud inputs [29].
@@ 3 Method
Figure 2: Overview of our method. Given input scene point cloud & robot pose, our method performs a series of diffusion steps to obtain denoised actions with SIM(3)-equivariance, i.e. when the inputs translate, rotate, and scale, the outputs are guaranteed to translate, rotate, and scale accordingly.
In this section, we describe the design of EquiBot. Our method builds upon recent advances that use a diffusion process to represent visuomotor policies [1]. Starting from that, we provide key insights for injecting equivariant architectures. By building an equivariant noise prediction network, we enforce each diffusion step to be equivariant by construction (see Figure 2), and because of the self-symmetry of the initial Gaussian noise, the overall framework outputs an equivariant distribution under stochasticity. Section C in supplemental materials provides our formal argument for this. The equivariant update in the diffusion process plays the role of letting the network “see” different input variations under transformations, which replaces the need for data augmentations and makes our framework data-efficient. Below, we introduce concepts related to equivariance and diffusion policy, then describe our method.
@@ 3.1 Preliminaries
Problem setup. We assume an imitation learning setup, where our method receives a demonstration dataset $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$, which consists of $N$ demonstration trajectories $\tau^{n}$. Each demonstration trajectory consists of sequences of observation-action pairs $({\mathbf{O}}_{t},{\mathbf{A}}_{t})$. The goal of the policy is to learn a mapping $\pi$ from observation ${\mathbf{O}}_{t}$ to either the next action ${\mathbf{A}}_{t}$ or the next set of actions ${\mathbf{A}}_{t:t+T_{p}}$, where $T_{p}$ is the prediction horizon. At evaluation time, the policy receives state ${\mathbf{O}}_{t}$ and predicts one or more actions to be executed in the environment. In this work, we assume an observation ${\mathbf{O}}_{t}=({\mathbf{X}}_{t},{\mathbf{S}}_{t})$ is composed of the scene point cloud ${\mathbf{X}}_{t}$ and robot proprioception state ${\mathbf{S}}_{t}$.
@@
${\mathbf{A}}_{t}^{k-1}=\mathbf{\alpha}({\mathbf{A}}_{t}^{k}-\gamma\epsilon_{\theta}({\mathbf{O}}_{t},{\mathbf{A}}_{t}^{k},k)+\mathcal{N}(0,\sigma^{2},I)).$
The policy outputs ${\mathbf{A}}_{t}^{0}$ as its inference output. In this work, we use the CNN-based Diffusion Policy variant specified in [1] as the starting point of our architecture design. Note that although our method uses point clouds as visual input format rather than RGB images, most architectural details of this policy can be adopted. The original CNN-based Diffusion Policy architecture uses a noise prediction network that takes observation ${\mathbf{O}}_{t}$, diffusion timestep $k$, and noisy action ${\mathbf{A}}_{t}$ as input, and predicts the gradient $\nabla{\mathbf{E}}({\mathbf{A}}_{t})$ for denoising the noisy action input. The network first uses an encoder to encode the visual observations. The encoded visual features and positional embeddings of the time step parameter are passed into FiLM layers [30] so that the encoded visual inputs are integrated into the network. Then, the policy network uses a convolutional U-net [31] to process the input noisy actions ${\mathbf{A}}_{t}$, the conditioned observations, and diffusion timestep $k$ to predict the output denoising gradients.
@@ 3.2 SIM(3)-Equivariant Diffusion Policy
To design a SIM(3)-equivariant model architecture, our approach is to modify each part of the CNN-based diffusion policy architecture [1] to make them individually equivariant architectures. First, we design our point cloud encoder to be SIM(3)-equivariant and additionally output a centroid vector $\Theta_{c}$, as well as a scalar $\Theta_{s}$ quantifying object scale. $\Theta_{c}$ and $\Theta_{s}$ are then used to scale the inputs to subsequent layers of the network, so that they are invariant to positions and scales. We then modify the FiLM layers, the convolutional U-net architecture, and other connecting layers to be SO(3)-equivariant (aka. equivariant to 3D rotations). Finally, before producing the output actions, we scale the relevant part of the action back using $\Theta_{c}$ and $\Theta_{s}$, so the output is SIM(3)-equivariant to the input observation. Please see supplementary materials (Figure 10) for a detailed method figure.
Encoder. We use a PointNet-based [32] encoder in this work. For SIM(3)-equivariance, we reuse the encoder $\Phi$ introduced in [21]. This encoder takes a point cloud ${\mathbf{X}}$ as input and outputs a latent code $\Theta=\Phi({\mathbf{X}})$, comprised of four components: $\Theta:=(\Theta_{R},\Theta_{\text{inv}},\Theta_{c},\Theta_{s})$, where $\Theta_{R}$ is a rotation equivariant latent representation, $\Theta_{\text{inv}}$ is an invariant latent representation, scalar $\Theta_{s}$ is the computed object scale, and vector $\Theta_{c}$ denotes the object centroid. For more details on this encoder, we refer to [21]. While [21] pre-trains the encoder using generated simulation data, we do not perform pre-training on the encoder and learn it from scratch. This eliminates the need to build task-specific simulation environments and collect custom pre-training data in these environments.
@@
To make 1D convolution layers SO(3)-equivariant, we treat vector channels of the layer inputs as batch dimensions and perform the original convolution operations. This simple change makes the convolution layer SO(3)-equivariant. We do not make any modifications to the upsampling layer, as it is naturally SO(3)-equivariant. To make the FiLM layer SO(3)-equivariant, we substitute vanilla linear layers with vectorized linear layers introduced in [33]. Formally, a FiLM layer is formulated as $\text{FiLM}({\mathbf{F}}|\gamma,\beta)=\gamma{\mathbf{F}}+\beta$, where $\gamma=f(\mathbf{x})$ and $\beta=h(\mathbf{x})$ are parameters predicted from learned functions used to modulate a neural network layer’s activations ${\mathbf{F}}$, and $\mathbf{x}$ is this neural network layer’s input. We replace non-equivariant layers $f$ and $h$ with ‘vector neuron’ layers [33], achieving rotation equivariance.
@@
Implementation Details. Data normalization can be important to the performance of diffusion models. Vanilla diffusion policy normalizes the observations and actions separately. We instead normalize all 3D-vector inputs (including positions and velocities) together due to our SIM(3)-equivariance assumptions. Implementation-wise, we take a subset of training data and compute the mean point cloud scale and mean action scale as $s_{\text{pc}}$ and $s_{\text{ac}}$. Then, all position- and velocity-related information is divided by $s_{\text{pc}}/s_{\text{ac}}$ at the start of the network forward pass and multiplied back before the output is returned. This normalization factor ensures that the diffusion process always works with actions with values within the $-1$ to $1$ range. We do not apply offsets to position and velocity information in this work. We normalize scalar information in the same way as the vanilla diffusion policy. Please see the supplementary materials (Section D) for further implementation details and model hyperparameters.

### v2

@@ Abstract
Building effective imitation learning methods that enable robots to learn from limited data and still generalize across diverse real-world environments is a long-standing problem in robot learning. We propose EquiBot, a robust, data-efficient, and generalizable approach for robot manipulation task learning. Our approach combines SIM(3)-equivariant neural network architectures with diffusion models. This ensures that our learned policies are invariant to changes in scale, rotation, and translation, enhancing their applicability to unseen environments while retaining the benefits of diffusion-based policy learning such as multi-modality and robustness. We show on a suite of 6 simulation tasks that our proposed method reduces the data requirements and improves generalization to novel scenarios. In the real world, with 10 variations of 6 mobile manipulation tasks, we show that our method can easily generalize to novel objects and scenes after learning from just 5 minutes of human demonstrations in each task. Website: https://equi-bot.github.io/
@@ 1 Introduction
The unpredictability and variability of real-world environments have posed significant challenges to the development of fully autonomous robotic agents. Existing visuomotor policies learned via imitation learning are effective in controlled settings [1, 2, 3, 4, 5, 6, 7], but they require substantial data and generalize poorly to unseen scenarios [8, 9, 10, 11, 12], limiting their practical deployment.
In this work, we introduce EquiBot, an equivariant policy learning architecture based on diffusion models [13]. The equivariant neural network guarantees that outputs scale, translate, and rotate with inputs, even if not fully trained [14], enabling generalization to unseen scenarios with changes in object visuals, sizes, and placements. This architecture also has data efficiency benefits, as it can infer actions for object placements and poses that are in distribution but insufficiently demonstrated, outperforming non-equivariant policies. Using a diffusion model-based policy architecture also offers robust learning performance, including support for idle actions and multi-modal behaviors [8]. Leveraging this design, EquiBot enables a wide range of household manipulation tasks using only a handful of single-view human demonstration videos. At test time, our method takes single-view, object-centric point clouds and robot proprioception as input and outputs sequences of robot end-effector actions that consist of 6D velocity commands and gripper open-close signals.
We evaluate our method in both simulation and real-world experiments. In simulation, we test on a suite of six tasks adapted from prior benchmarks [8, 15, 16], showing that our method is more data-efficient and generalizes better to unseen scenarios than prior works. In real robot experiments, we quantitatively test on 6 real-world mobile manipulation tasks in home settings, including pushing a chair towards a desk, closing a laundry machine door, folding towels, making a bed, and closing a suitcase. Our method successfully performs these tasks with unseen objects and scenes from just 5 minutes of human demonstrations, outperforming baselines that rely on augmentations for generalization or do not use a diffusion process to predict actions.
@@ 2 Related Work
Data-efficient imitation learning. Recent imitation learning approaches for imitation learning assume large quantities of data to be available for learning manipulation policies [4, 8]. To reduce the amount of data that policy training requires per task, some work [2, 17, 18, 19] formulate a one-shot or few-shot imitation learning setup where the policy is trained on multi-task demonstration data and can then output actions in a novel task after seeing one or more demonstrations as well as the current state. However, this approach requires the availability of multi-task data in a task domain. Some other works [20] achieve imitation learning from small demonstration datasets with sampling-based optimization methods like Bayesian Optimization, but these methods are often limited to small action spaces and open-loop settings. In contrast to prior works, we show that embedding equivariance into the policy architecture can effectively improve the data efficiency of the imitation learning algorithm, allowing robust policies to be learned with just a handful of demonstrations.
@@
Equivariance in robot manipulation. To help learned robot policies generalize to unseen environments and object placements, prior works [21, 22, 23] use data augmentation to improve cross-domain transfer of the learned policy. However, these methods increase training time significantly and do not guarantee generalization to visual appearances, object scales, and poses that are unseen in the training data distribution even after the augmentation. In contrast, utilizing equivariant representations in policy learning allows the learned policies to generalize to objects and initial conditions not previously seen at training time. Prior works have explored the use of equivariance in robot manipulation in several different setups [24, 25, 26, 27, 28, 29], but most of them either focus on only simple pick-and-place like tasks, do not support closed-loop policies, or do not support scale equivariance. In our prior work [16], we developed a SIM(3)-equivariant visuomotor policy learning method that can go beyond pick-and-place tasks with deformable and articulated objects. However, EquivAct cannot handle multi-modal training data because of its deterministic architecture. EquivAct also lacks the necessary architectural design that prevent it from predicting multiple actions into the future, which is crucial for producing temporally consistent actions [8]. Compared to prior works, our method trains stably, handles multi-modal data with diffusion, and generalizes to unseen object appearance, initial states, scales, and poses with SIM(3)-equivariance.
@@
Equivariant diffusion architectures. Prior works have integrated equivariance in diffusion models in various non-robotics domains [30, 31, 32, 33]. Some works have attempted to integrate equivariant architectures in diffusion models for robotics [34, 35, 36]. Diffusion-EDFs [34] produces target end-effector poses and can only solve pick-and-place tasks. EquiDiff [35] can only handle SO(2)-equivariance with simple 2D trajectories. EDGI [36] assumes ground-truth scene states as input and cannot handle complex visual observations. Compared to prior works, our work proposes an equivariant diffusion policy for closed-loop 3D manipulation tasks with point cloud observations.
@@ 3 Method
In this section, we describe the design of EquiBot. Our method builds upon recent advances that use a diffusion process to represent visuomotor policies [8]. Starting from that, we provide key insights for injecting equivariant architectures. By building an equivariant noise prediction network, we enforce each diffusion step to be equivariant by construction (see Figure 2), and because of the self-symmetry of the initial Gaussian noise, the overall framework outputs an equivariant distribution under stochasticity. Section 3.2 provides our formal argument for this. The equivariant update in the diffusion process plays the role of letting the network “see” different input variations under transformations, which replaces the need for data augmentations and makes our framework data-efficient. Below, we introduce concepts related to equivariance and diffusion policy, then describe our method.
@@ 3.1 Preliminaries
Figure 2: Method overview. Given input scene point cloud & robot pose, our method performs a series of diffusion steps to obtain denoised actions with SIM(3)-equivariance, i.e. when the inputs translate, rotate, and scale, the outputs are guaranteed to translate, rotate, and scale accordingly.
Problem setup. We assume an imitation learning setup, where our method receives a demonstration dataset $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$, which consists of $N$ demonstration trajectories $\tau^{n}$. Each demonstration trajectory consists of sequences of observation-action pairs $({\mathbf{O}}_{t},{\mathbf{A}}_{t})$. The goal of the policy is to learn a mapping $\pi$ from past observations ${\mathbf{O}}_{t-T_{o}:t}$ to either the next action ${\mathbf{A}}_{t}$ or the next set of actions ${\mathbf{A}}_{t:t+T_{p}}$, where $T_{o}$ and $T_{p}$ are the observation and prediction horizons. At evaluation time, the policy receives state ${\mathbf{O}}_{t}$ and predicts the next one or more actions to be executed in the environment. In this work, we assume the observation ${\mathbf{O}}_{t}=({\mathbf{X}}_{t},{\mathbf{S}}_{t})$ is composed of the scene point cloud ${\mathbf{X}}_{t}$ and robot proprioception ${\mathbf{S}}_{t}$.
@@
${\mathbf{A}}_{t}^{k-1}=\mathbf{\alpha}_{k}({\mathbf{A}}_{t}^{k}-\gamma_{k}\mathbf{\epsilon}_{\theta}({\mathbf{O}}_{t},{\mathbf{A}}_{t}^{k},k)+\sigma_{k}\mathcal{N}(0,{\mathbf{I}})),$ (1)
where $\epsilon_{\theta}$ is a denoising network, $\mathcal{N}(0,{\mathbf{I}})$ is Gaussian noise, and $\alpha_{k},\gamma_{k},\sigma_{k}$ are functions of $k$ set by a noise scheduler. The policy outputs ${\mathbf{A}}_{t}^{0}$ as its inference output. In this work, we use the CNN-based Diffusion Policy variant specified in [8] as the starting point of our architecture design. The original CNN-based Diffusion Policy architecture uses a noise prediction network that takes observation ${\mathbf{O}}_{t}$, diffusion iteration $k$, and noisy action ${\mathbf{A}}_{t}$ as input, and predicts the gradient $\nabla{\mathbf{E}}({\mathbf{A}}_{t})$ for denoising ${\mathbf{A}}_{t}$. The network first uses an encoder to encode the visual observations. The encoded visual features and positional embeddings of the diffusion iteration parameter are passed into FiLM layers [37] so that the encoded visual inputs are integrated into the network. Then, the policy network uses a convolutional U-net [38] to process the input noisy actions ${\mathbf{A}}_{t}$, the conditioned observations, and diffusion iteration $k$ to predict the output denoising gradients.
@@
3.2 Equivariant Distributions and Diffusion
We make a diffusion process equivariant by making the architecture for performing each diffusion step equivariant. In this section, we show that with equivariance in the per-step diffusion update, the final output action distribution is also equivariant under stochasticity. We mainly discuss equivariance to SO(3)-rotations in the diffusion process. Equivariance to translations and scaling (to get SIM(3)-equivariant architectures) is achieved via canonicalization before the diffusion process.
Proposition 1.
Let $p(x^{K}|c)$ be an SO(3)-equivariant density function conditioned on $c$, i.e. $\forall{\mathbf{R}}\in SO(3),p(x^{K}|c)=p({\mathbf{R}}x^{K}|{\mathbf{R}}c)$. If the Markov transitions $p(x^{k-1}|x^{k},c)$ are SO(3)-equivariant for all $k$, i.e. $p(x^{k-1}|x^{k},c)=p({\mathbf{R}}x^{k-1}|{\mathbf{R}}x^{k},{\mathbf{R}}c)$, then the density $p(x^{0}|c)=\int p(x^{T}|c)\Pi_{k=1}^{K}p(x^{k-1}|x^{k},c)$ is also SO(3)-equivariant.
Please view the complete proof in Section D of the supplementary materials. In our case, observation ${\mathbf{O}}_{t}$ conditions the diffusion process of the actions ${\mathbf{A}}_{t}^{0:K}$. The prior distribution $p({\mathbf{A}}_{t}^{K}|{\mathbf{O}}_{t})$ is a standard Gaussian distribution equivariant to SO(3) transformations. The transition probabilities $p({\mathbf{A}}_{t}^{k-1}|{\mathbf{A}}_{t}^{k},{\mathbf{O}}_{t})$ are predictions by an equivariant network, and thus are equivariant to SO(3)-rotations. Therefore, the final output action $p({\mathbf{A}}_{t}^{0}|{\mathbf{O}}_{t})$ is also SO(3)-equivariant.
@@ 3.3 SIM(3)-Equivariant Diffusion Policy
To design a SIM(3)-equivariant model architecture, our approach is to modify each part of the CNN-based Diffusion Policy architecture [8] to make them individually equivariant architectures. First, we design our point cloud encoder to be SIM(3)-equivariant and additionally output a centroid vector $\Theta_{c}$ as well as a scalar $\Theta_{s}$ quantifying object scale. $\Theta_{c}$ and $\Theta_{s}$ are then used to scale the inputs to subsequent layers of the network so that they are invariant to positions and scales. We then modify the FiLM layers, the convolutional U-net architecture, and other connecting layers to be SO(3)-equivariant. Finally, before producing the output actions, we scale relevant parts of the action back using $\Theta_{c}$ and $\Theta_{s}$ so the output is SIM(3)-equivariant to the input observation. Please see supplementary materials (Figure 13) for a detailed method figure.
Encoder. We use a PointNet-based [39] encoder in this work. For SIM(3)-equivariance, we reuse the encoder $\Phi$ introduced in [40] and [16]. This encoder takes a point cloud ${\mathbf{X}}$ as input and outputs a latent code $\Theta=\Phi({\mathbf{X}})$, comprised of four components: $\Theta:=(\Theta_{R},\Theta_{\text{inv}},\Theta_{c},\Theta_{s})$, where $\Theta_{R}$ is a rotation equivariant latent representation, $\Theta_{\text{inv}}$ is an invariant latent representation, scalar $\Theta_{s}$ is the computed object scale, and vector $\Theta_{c}$ denotes the object centroid. For more details on this encoder, we refer to [16]. While [16] pre-trains the encoder using generated simulation data, we do not perform pre-training on the encoder and learn it from scratch. This eliminates the need to build task-specific simulation environments and collect custom pre-training data in these environments.
@@
To make 1D convolution layers SO(3)-equivariant, we treat vector channels of the layer inputs as batch dimensions and perform the original convolution operations. This simple change makes the convolution layer SO(3)-equivariant. We do not make any modifications to the upsampling layer, as it is naturally SO(3)-equivariant. To make the FiLM layer SO(3)-equivariant, we substitute vanilla linear layers with vectorized linear layers introduced in [14]. More formally, a FiLM layer is formulated as $\text{FiLM}({\mathbf{F}}|\eta,\beta)=\eta{\mathbf{F}}+\beta$, where $\eta=f(\mathbf{x})$ and $\beta=h(\mathbf{x})$ are parameters predicted from learned functions used to modulate a neural network layer’s activations ${\mathbf{F}}$, and $\mathbf{x}$ is this neural network layer’s input. We replace non-equivariant layers $f$ and $h$ with ‘vector neuron’ layers [14], achieving rotation equivariance.
@@
Please refer to Section E in supplementary materials for implementation details.